Using the Chain Rule, find the 𝑑𝑦/𝑑π‘₯ and express the final answer in term of x.

𝑦 = 2𝑒/π‘’Β²βˆ’1 , u=xΒ²


1
Expert's answer
2022-03-03T17:49:51-0500

Solution


Given that


"y=\frac{2u}{u^{2}-1} ... (1)\\
\\
u=x^{2} ... (2)"


From (1),


"{\color{Blue} y=\frac{2u}{u^{2}-1}}\\
\frac{dy}{du}=\frac{(u^{2}-1)\frac{d}{du}(2u)-(2u)\frac{d}{du}(u^{2}-1)}{(u^{2}-1)^{2}}\\
\frac{dy}{du}=\frac{(u^{2}-1)(2)-(2u)(2u)}{(u^{2}-1)^{2}}\\
\frac{dy}{du}=\frac{(2u^{2}-2)-(4u^{2})}{(u^{2}-1)^{2}}\\
\frac{dy}{du}=\frac{(-2u^{2}-2)}{(u^{2}-1)^{2}}\\
\frac{dy}{du}=\frac{-2(u^{2}+1)}{(u^{2}-1)^{2}} ... (3)\\"


Now from (2)

"{\color{Blue}u=x^{2}}\\
\frac{du}{dx}=\frac{d}{dx}(x^{2})\\
\frac{du}{dx}=2x ... (4)"


According to the chain rule

"\frac{dy}{dx}=\frac{dy}{du}\times \frac{du}{dx}"



Using (3) and (4)


"\frac{dy}{dx}=\frac{dy}{du}\times \frac{du}{dx}\\
\\
\\

\frac{dy}{dx}=\frac{-2(u^{2}+1)}{(u^{2}-1)^{2}}\times (2x)\\
\\
\\

\frac{dy}{dx}=\frac{-4x(u^{2}+1)}{(u^{2}-1)^{2}}\\"


Hence according to the chain rule,

"{\color{Blue}\frac{dy}{dx}=\frac{-4x(u^{2}+1)}{(u^{2}-1)^{2}}}\\"